# MAINS: A Magnetic Field Aided Inertial Navigation System for Indoor Positioning

### v2

@@ II-C Tranforming models between body frames
Next, substituting the generic magnetic model in (II-C) with the proposed polynomial model (10) yields the equality
@@ II-E Complete System
respectively. Combining the models in (II-A), (19), and (21) gives the state-space model
@@ III State Estimation
The quaternion $q_{k}$ in the state vector does not belong to Euclidean space. Therefore, standard nonlinear filter algorithms, such as the extended Kalman filter and unscented Kalman filter, cannot be applied to the state-space model in (II-E) without appropriate modifications. The error state Kalman filter (ESKF) presented in [25] circumvents the problem by the use of “error quaternion”, which is a small perturbation around the “estimated quaternion” expressed in $\mathbb{R}^{3}$. With reference to Fig. 4, the algorithm works by propagating an estimated state $\hat{x}_{k}$ via the state transition model in (23a) and using a complementary Kalman filter to estimate the error state $\delta x_{k}$, which is then used to correct the estimated state $\hat{x}_{k}$. Since the ESKF is a standard algorithm, only the error propagation model that is unique to the state-space model in (II-E), will here be derived.
@@ III-B Inertial Error State Dynamics
$G^{\text{\tiny{ins}}}_{k}$ $=\begin{bmatrix}0&0&0&0\\ \hat{R}_{b_{k}}^{n}T_{s}&0&0&0\\ 0&I_{3}T_{s}&0&0\\ 0&0&I_{3}\sqrt{T_{s}}&0\\ 0&0&0&I_{3}\sqrt{T_{s}}\end{bmatrix}.$ (27h)
@@ III-C Magnetic Field Subsystem Error State Dynamics
$F^{\theta}_{k}$ $=$ $A^{\dagger}\begin{bmatrix}B(\hat{\psi}_{k})&J_{1}&J_{2}\end{bmatrix}C(\hat{x}_{k}),$ (32c)
$G^{\theta}_{k}=\begin{bmatrix}0\quad-A^{\dagger}J_{2}T_{s}\quad 0\quad 0\quad I_{\kappa}\end{bmatrix}.$ (32d)
@@
$J_{1}$ $=$ $\frac{d}{d\Delta p_{k}}\bigl{(}B(\hat{\psi}_{k})\theta_{k}\bigr{)},$ (32e)
$J_{2}$ $=$ $\frac{d}{d\Delta\phi_{k}}\bigl{(}B(\hat{\psi}_{k})\theta_{k}\bigr{)},$ (32f)
@@
$\!C(\hat{x}_{k})\!=\!\begin{bmatrix}0&\!0&\!0&\!0&\!0&\!I_{\kappa}\\ 0&\!\hat{R}^{b_{k}}_{n}T_{s}&\![\eta(\hat{R}^{b_{k}}_{n},\hat{v}^{n}_{k})]_{\times}&\!0&\!0&\!0\\ 0&\!0&\!0&\!0&\!-I_{3}T_{s}&\!0\end{bmatrix}.$ (32g)
Combining (III-B) and (III-C) gives the complete state-space model of the error state, i.e.,
@@
$F_{k}=\begin{bmatrix}F^{\text{\tiny{ins}}}_{k}&0\\ \hfil F^{\theta}_{k}\hfil\end{bmatrix},\quad G_{k}=\begin{bmatrix}G^{\text{\tiny{ins}}}_{k}&0\\ \hfil G^{\theta}_{k}\hfil\end{bmatrix},$ (33c)
$H_{\delta x}=\begin{bmatrix}0_{3\times 15}&\Phi(r_{m_{1}})\\ \vdots&\vdots\\ 0_{3\times 15}&\Phi(r_{m_{N}})\\ \end{bmatrix}.$ (33g)
@@
The Kalman filter can be applied to (III-C) to estimate the error state, which is then used to correct the estimated state. The complete description of the proposed ESKF is summarized in Algorithm 1.

### v3

@@ II-C Tranforming models between body frames
Next, substituting the generic magnetic model in (14) with the proposed polynomial model (10) yields the equality
@@ II-E Complete System
respectively. Combining the models in (2), (19), and (21) gives the state-space model
@@ III State Estimation
The quaternion $q_{k}$ in the state vector does not belong to Euclidean space. Therefore, standard nonlinear filter algorithms, such as the extended Kalman filter and unscented Kalman filter, cannot be applied to the state-space model in (23) without appropriate modifications. The error state Kalman filter (ESKF) presented in [25] circumvents the problem by the use of “error quaternion”, which is a small perturbation around the “estimated quaternion” expressed in $\mathbb{R}^{3}$. With reference to Fig. 4, the algorithm works by propagating an estimated state $\hat{x}_{k}$ via the state transition model in (23a) and using a complementary Kalman filter to estimate the error state $\delta x_{k}$, which is then used to correct the estimated state $\hat{x}_{k}$. Since the ESKF is a standard algorithm, only the error propagation model that is unique to the state-space model in (23), will here be derived.
@@ III-B Inertial Error State Dynamics
$G^{\text{\tiny{ins}}}_{k}$ $=\begin{bmatrix}0&0&0&0\\ \hat{R}_{b_{k}}^{n}T_{s}&0&0&0\\ 0&I_{3}T_{s}&0&0\\ 0&0&I_{3}\sqrt{T_{s}}&0\\ 0&0&0&I_{3}\sqrt{T_{s}}\end{bmatrix}.$ (27c)
@@ III-C Magnetic Field Subsystem Error State Dynamics
$F^{\theta}_{k}$ $=$ $A^{\dagger}\begin{bmatrix}B(\hat{\psi}_{k})&J_{1}&J_{2}\end{bmatrix}C(\hat{x}_{k}),$ (32b)
$G^{\theta}_{k}=\begin{bmatrix}0\quad-A^{\dagger}J_{2}T_{s}\quad 0\quad 0\quad I_{\kappa}\end{bmatrix}.$ (32c)
@@
$J_{1}$ $=$ $\frac{d}{d\Delta p_{k}}\bigl{(}B(\hat{\psi}_{k})\theta_{k}\bigr{)},$ (32d)
$J_{2}$ $=$ $\frac{d}{d\Delta\phi_{k}}\bigl{(}B(\hat{\psi}_{k})\theta_{k}\bigr{)},$ (32e)
@@
$\!C(\hat{x}_{k})\!=\!\begin{bmatrix}0&\!0&\!0&\!0&\!0&\!I_{\kappa}\\ 0&\!\hat{R}^{b_{k}}_{n}T_{s}&\![\eta(\hat{R}^{b_{k}}_{n},\hat{v}^{n}_{k})]_{\times}&\!0&\!0&\!0\\ 0&\!0&\!0&\!0&\!-I_{3}T_{s}&\!0\end{bmatrix}.$ (32f)
Combining (27) and (32) gives the complete state-space model of the error state, i.e.,
@@
[…]
$H_{\delta x}=\begin{bmatrix}0_{3\times 15}&\Phi(r_{m_{1}})\\ \vdots&\vdots\\ 0_{3\times 15}&\Phi(r_{m_{N}})\\ \end{bmatrix}.$ (33d)
@@
The Kalman filter can be applied to (33) to estimate the error state, which is then used to correct the estimated state. The complete description of the proposed ESKF is summarized in Algorithm 1.